A long jumper reaches a horizontal velocity of 6ms just prior to leaving the ground. If she jumps a distance of 6 m from her starting point, modeling what was the highest point in her trajectory? Can you please show all work? I have tried this problem repeatedly and cannot seem to get to the answer given as 1.23 meters
Answer #1

Range , which is horizontal distance = v2 sin(2\Theta)/g = 6 m

2vsin\Theta vcos\Theta/g = 6

v cos\Theta = 6 m/s (given, horizontal velocity)

2v sin\Theta * 6 /9.8 = 6

v sin\Theta = 4.9 m/s

Maximum height = (vsin\Theta)2/2g = 4.9*4.9/19.6 = 1.225 m which is approximately 1.23 meters